OpenStudy (anonymous):

How to solve for the equation 3x +5

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

is that all the question gives you?

OpenStudy (anonymous):

3X +5 =

OpenStudy (anonymous):

x=-5/3

OpenStudy (anonymous):

?? if u r solving for x

OpenStudy (anonymous):

ok so ur solving for x thats right

OpenStudy (radar):

That would only be true, if..... 3x+5=0

OpenStudy (radar):

As presented there is no solution.

OpenStudy (anonymous):

no .if f(x) =3x +2, find each value. f(x +1) =3x +5 f(x +1) = 3x +5 do i leave my answer as it is (( f(x +1) =3X +5 ))

OpenStudy (anonymous):

srry the first one should be 3(x+1) +2

OpenStudy (radar):

What is your problem?

OpenStudy (radar):

You keep changing the conditions!

OpenStudy (radar):

First you say it is an equation (which it wasn't as there was no = sign. Then you put an equal sign without any value. Then you say it is a function... Again what is your problem?

OpenStudy (radar):

Lets assume that f(x)= 3x+2 as you state. then f(x+1)=3(x+1)+2=3x+3+2=3x+5 now you indicate: 3x+5=3x+5 seems reasonable to me.

OpenStudy (anonymous):

Thanks to everyone who helped out with my work